OpenStudy (anonymous):

1024=x^3 how do i simplify it?

OpenStudy (anonymous):

do know about cube root

OpenStudy (anonymous):

1024 = x^3 x = (1024)^1\3

OpenStudy (anonymous):

so 1024 divide by 3?

OpenStudy (anonymous):

no no cube root of 1024

OpenStudy (anonymous):

ok i get it know

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

What is the cube root of 1024? @savvers

OpenStudy (anonymous):

(1024)^(1/3) = 8*(2)^(1/3) = 8*cubrt(2) cubrt(a*b) = cubrt(a)*cubrt(b) Given, 1024 = 8*8*8*2 cubrt(1024) = cubrt(8*8*8*2) = cubrt(8)*cubrt(8)*cubrt(8)*cubrt(2) cubrt(8) = 2, so 2*2*2*cubrt(2) or 8*cubrt(2) The simplest form is x = 8*cubrt(2)

OpenStudy (anonymous):

@savvers @uzumakhi